An airplane is flying at Mach 1.63 at an altitude of 7,200.00 meters, where the speed of sound is v = 314.33 m/s. How far away from a stationary observer will the plane be when the observer hears the sonic boom? (Enter the total distance in m.)
Answer #1

time taken by the sound to reach ground observer

t = 7200/ 314.33 = 22.906 seconds

distance travelled by the plane till this time horizontally

x = 1.63 * 314.33 * 22.906= 11736.07 m

now, net displacement

d^2 = x^2 + y^2

d^2 = 11736.07^2 + 7200^2

d = 13768.636 m

=========

Comment in case any doubt, will reply for sure.. Goodluck